Four dice are rolled and sum of numbers which appeared is noted as x.find the probability that x=13
Answer by rothauserc(4718)

The sample space for rolling 4 6-sided dice is 6^4 = 1296
:
The number of ways you can get a sum of 13 is 140
:
Probability (rolling a 13 with 4 6-sided dice = 140/1296 = 0.108
:
The intuitive way to calculate the number of ways we can get a sum of 13 with 4 die follows
:
each die has a minimum of 1 dot and a maximum of 6 dots
:
force each die to have 1 dot, this leaves us 9 dots to distribute to the 4 die
:
If there wasn’t an upper limit for the points to assign to the dice we’d be done. We would apply the combination with repetition formula to find all the ways
:
((9C4)) = ((4+9-1)C9) = 12C9 = 12!/(9!(12-9)!) = 220 ways
:
BUT, we have to account for the outlawed ways, that is, dice with a number of points ≥ 6, then
:
number of allowed ways = number of unrestrained ways – number of outlawed ways
:
consider 1 outlawed die with a 6, then that leaves 3 dots to distribute to the 4 die
:
((4C3)) = ((4+3-1)C3) = (6C3) = 6!/(3!(6-3)!) = 20
:
there are 4 die, so we need to multiply by 4, 4 * 20 = 80
:
number of allowed ways = 220 - 80 = 140
: